### v1

@@ 3.2. Dislocated Jacobi operators
In order to study the right–half Jacobi operator $H^{\Sigma}_{R}(t)$, we study yet another dislocated model. The idea to study this model comes from Hempel and Kohlmann in [21, 22], and was used in [16, 18] to compute some spectral flows for continuous Schrödinger operators acting on a one-dimensional channel. In this setting, the dislocated model is modelled by a Schrödinger operator of the form $H_{t}:=-\partial_{xx}^{2}+V_{t}(x)$ with the dislocated potential
@@
where $V$ is $1$–periodic. The map $t\mapsto H_{t}$ is norm–resolvent continuous and $1$-periodic, and we recover periodic models at the endpoints. As mentioned in the introduction, a difficulty with the discrete model is that it is not possible to create such a continuous shift. Our workaround is to introduce a quasi-periodic dislocated model $\widetilde{H}^{\Sigma}(t)$.
@@
The operator $\widetilde{H}^{\Sigma}(0)$ is exactly the periodic Jacobi operator $H$, and its spectrum is $\sigma_{\rm{bulk}}$. For the operator $\widetilde{H}^{\Sigma}(1)$, we note the following. Let us write ={0}($\mathbb{Z}=\{0\}\cup(\mathbb{Z}\setminus\{0\})$, and $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})=\mathbb{C}^{N}\oplus\ell^{2}(\mathbb{Z}\setminus\{0\},\mathbb{C}^{N})$. With respect to this decomposition, we write
@@
$\widetilde{H}^{\Sigma}(1)=\Sigma\,\widetilde{\oplus}\,\widetilde{H}^{\Sigma}(0)$ (13)
@@ 3.3. Computation of the spectral flow for the dislocated model
This result was first proved in [21, 22] in the case of continuous Schrödinger operators. The proof presented here is similar, but more complete, and uses solely tools from spectral theory.
@@ 3.3.1. The finite periodic dislocated model
At $t=1$, apart from the presence of matrix $\Sigma$ in the middle, we also recognize a periodic model, and we have a relation of the form (compare with (13))
$\widetilde{H}_{\ell}^{\Sigma}(1)=\Sigma\,\widetilde{\oplus}\,H_{\ell-1},\qquad\text{hence}\qquad\sigma\left(\widetilde{H}_{\ell}^{\Sigma}(1)\right)=\sigma\left(H_{\ell-1}\right)\cup\{\Sigma\},$
@@ 3.3.2. From the finite dislocated model to the infinite dislocated model
In the previous section, we proved that the spectral flow of the supercell dislocated model $\widetilde{H}^{\Sigma}_{\ell}$ is $-{\mathcal{N}}(E)$ for all $\ell\geq 5$. In this section, we justify that one may take the limit $\ell\to\infty$ and deduce that spectral flow of is also $-{\mathcal{N}}(E)$ for the full dislocated model $\widetilde{H}^{\Sigma}$. Throughout this subsection, we drop the superscript $\Sigma$, and the tilde notation for clarity: we will simply write $H(t)$ and $H_{\ell}(t)$ for the dislocated models acting on $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})$ and $\ell^{2}(\mathbb{Z},\Omega_{\ell})$ respectively. Recall that at $t=0$, the operators $H(0)$ and $H_{\ell}(0)$ are periodic, and that
@@
$\chi(x)=1\quad\forall x\in[-1/4,1/4],\qquad\chi(x)=0\quad\forall x\in\mathbb{R}\setminus(-1/3,1/3)$
and define $\chi_{\ell}$ as the multiplication operator by $\chi(\cdot/\ell)$. Note that this function is supported in $[-\ell/3,\ell/3]\subset\Omega_{\ell}$. For this reason, we abuse notation and use $\chi_{\ell}$ for multiplication by $\chi_{\ell}$ from and to ${\mathcal{H}}_{\ell}$ and ${\mathcal{H}}$ interchangeably. In other words, we identify the operator $\chi_{\ell}$ in ${\mathcal{H}}$ with $i_{\ell}^{*}\chi_{\ell}i_{\ell}$ in ${\mathcal{H}}_{\ell}$, and $i_{\ell}^{*}\chi_{\ell}$, $\chi_{\ell}i_{\ell}$ between ${\mathcal{H}}$ and ${\mathcal{H}}_{\ell}$. We will use the following identities throughout this section.
@@
    $H(t)\chi_{\ell}=i_{\ell}H_{\ell}(t)\chi_{\ell}$ and $\chi_{\ell}H(t)=\chi_{\ell}H_{\ell}(t)i_{\ell}^{*}$,
@@
    $H(t)(1-\chi_{\ell})=H(0)(1-\chi_{\ell})$ and $H_{\ell}(t)(1-\chi_{\ell})=H_{\ell}(0)(1-\chi_{\ell})$,
@@
    as operators from ${\mathcal{H}}$ to ${\mathcal{H}}_{\ell}$, we have
    $H_{\ell}(t)\chi_{\ell}=\chi_{\ell}H(t)+R_{\ell},\qquad\text{ with $R_{\ell}$ independent of $t$ satisfying}\quad\|R_{\ell}\|_{\rm{op}}\leq\frac{C}{\ell}.$ (14)
@@
For the first point, we note that for all $\Psi\in{\mathcal{H}}$, the function $\chi_{\ell}\Psi$ has support in $[-\ell/3,\ell/3]$. Since $H(t)$ is a Jacobi operator, the function $H(t)\chi_{\ell}\Psi$ has support in $[-\ell/3-1,\ell/3+1]$, and can be identified with a function in ${\mathcal{H}}_{\ell}$ by applying $i_{\ell}^{*}$, whenever $\ell\geq 12$. The second relation is the adjoint of the first.
For the second point, we note that $H(t)=H(0)+M(t)$, where $M(t)$ is a finite rank operator with nonzero matrix elements only in the entries $\{-1,0,1\}$. So $M(t)(1-\chi_{\ell})=0$ whenever $\ell\geq 12$, since $\chi_{\ell}(i)=1$ for $|i|\leq\ell/4$. The result follows.
For the third point, we first recall that $H(t)=H(0)+M(t)$ and $H_{\ell}(t)=H_{\ell}(0)+i_{\ell}M(t)i_{\ell}^{*}$. We already proved that $M_{\ell}(t)\chi_{\ell}=\chi_{\ell}M(t)$ for all $\ell\geq 12$. Next, we find that
$\left[H_{\ell}(0)\chi_{\ell}-\chi_{\ell}H(0)\right]_{ij}=\begin{cases}a\ (\chi_{\ell}(n+1)-\chi(n))&\quad\text{if}\quad(i,j)=(n,n+1)\\ a^{*}(\chi_{\ell}(n-1)-\chi(n))&\quad\text{if}\quad(i,j)=(n,n-1)\\ 0&\quad\text{else}\end{cases}$
@@
$R_{\ell}=H_{\ell}(t)\chi_{\ell}-\chi_{\ell}H(t)=H_{\ell}(0)\chi_{\ell}-\chi_{\ell}H(0)\quad\text{satisfies}\quad\|R_{\ell}\|\leq 2\|a\|_{\rm op}\frac{C}{\ell}.\qed$
First, we prove that points in the resolvent set of $H(t)$ are in the resolvent set of $H_{\ell}(t)$ for sufficiently large values of $\ell$.
@@
Since the essential spectrum of $H(t)$ coincides with $\sigma_{\rm{bulk}}$, we find that $\operatorname{dist}(\lambda,\sigma_{\rm{bulk}})\geq\varepsilon$. Recall that $\sigma(H_{\ell}(0))\subset\sigma_{\rm{bulk}}$, so in particular $H_{\ell}(0)-\lambda$ is invertible with $\|(H_{\ell}(0)-\lambda)^{-1}\|_{\rm op}\leq\varepsilon^{-1}$. The same bound holds for $(H(t)-\lambda)^{-1}$. We combine both resolvents to construct an approximate inverse for $H_{\ell}(t)-\lambda$. Indeed, for any $\ell\geq 12$,
$(H_{\ell}(t)-\lambda)\left\{\chi_{\ell}(H(t)-\lambda)^{-1}i_{\ell}+(1-\chi_{\ell})(H_{\ell}(0)-\lambda)^{-1}\right\}$
$\quad=\chi_{\ell}i_{\ell}+R_{\ell}(H(t)-\lambda)^{-1}i_{\ell}+(1-\chi_{\ell})-R_{\ell}(H_{\ell}(0)-\lambda)^{-1}$
$\quad=1+R_{\ell}\left\{(H(t)-\lambda)^{-1}i_{\ell}-(H_{\ell}(0)-\lambda)^{-1}\right\}:=1+\mathcal{R}_{\ell}(t).$
@@
$\left\|\mathcal{R}_{\ell}(t)\right\|_{\rm{op}}\leq\frac{2C}{\varepsilon\ell},$
@@
$\|(H_{\ell}(t)-\lambda)^{-1}\|_{\rm{op}}$ $\leq\left\|\chi_{L}(H(t)-\lambda)^{-1}i_{L}+(1-\chi_{L})(H_{\ell}(0)-\lambda)^{-1}\right\|_{\rm{op}}\,\|(1+\mathcal{R}_{\ell}(t))^{-1}\|_{\rm{op}}$
@@
$\forall z\in{\mathscr{C}},\quad\|(z-H_{\ell}(t))^{-1}\|_{\rm op}\leq\frac{4}{\varepsilon},\quad\text{and}\quad\|(z-H_{\ell}(t))^{-1}\|_{\rm op}\leq\frac{1}{\varepsilon}.$
We use the localization function $\chi_{\ell}$ and $\phi_{\ell}:=\sqrt{1-\chi_{\ell}^{2}}$, such that $\chi_{\ell}^{2}+\phi_{\ell}^{2}=1$. The support of $\chi_{\ell}$ is in $[-\ell/3,\ell/3]$, the support of $\phi_{\ell}$ is the complement of $(-\ell/4,\ell/4)$, and there is a constant such that $\|H_{\ell}(t)\chi_{\ell}-\chi_{\ell}H(t)\|_{{\rm{op}}}\leq C/\ell$.
For $z$ on ${\mathscr{C}}$, we obtain from Lemma 3.6, (as operators from ${\mathcal{H}}$ to ${\mathcal{H}}_{\ell}$)
$(H_{\ell}(t)-z)\chi_{\ell}=\chi_{\ell}(H(t)-z)+R_{\ell}.$
Multiplying this identity by $(H_{\ell}(t)-z)^{-1}$ on the left and $(H(t)-z)^{-1}$ on the right gives
$\chi_{\ell}(H(t)-z)^{-1}=(H_{\ell}(t)-z)^{-1}\chi_{\ell}+{\mathcal{R}}_{\ell}(z),$ (15)
with $\|{\mathcal{R}}_{\ell}(z)\|_{{\rm{op}}}\leq\frac{4}{\varepsilon^{2}\ell}$. We also have $(H_{\ell}(t)-z)^{-1}\phi_{\ell}^{2}=(H_{\ell}(0)-z)^{-1}\phi_{\ell}^{2}$, independent of $t$. Hence, we obtain
@@
$P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t))=\chi_{\ell}P_{(\lambda_{1},\lambda_{2})}(H(t))\chi_{\ell}+P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(0))\phi_{\ell}^{2}-\frac{1}{2{\mathrm{i}}\pi}\int_{\mathscr{C}}{\mathcal{R}}_{\ell}(z)\chi_{\ell}{\mathrm{d}}z.$
Since $(\lambda_{1},\lambda_{2})$ lies in a gap of $\sigma_{\rm{bulk}}$, the second term is actually zero. The last term can be bounded by $2|{\mathscr{C}}|/(\pi\varepsilon^{2}\ell)$. Thus, we can find $\ell_{0}$ sufficiently large such that for all $\ell\geq\ell_{0}$, we have
$\left\|P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t))-\chi_{\ell}P_{(\lambda_{1},\lambda_{2})}(H(t))\chi_{\ell}\right\|_{\rm{op}}<1.$
@@
For the opposite inequality, we return to (15) and write this time
$(H(t)-z)^{-1}$ $=\chi_{\ell}(H_{\ell}(t)-z)^{-1}\chi_{\ell}+\chi_{\ell}{\mathcal{R}}_{\ell}(z)+\phi_{\ell}^{2}(H(0)-z)^{-1}.$
Applying as before the Cauchy formula (the third term vanishes) and Lemma 3.9 gives
$\mathrm{rank}(P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t)))\geq\mathrm{rank}\left(\chi_{\ell}P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t))\chi_{\ell}\right)\geq\mathrm{rank}\left(P_{(\lambda_{1},\lambda_{2})}(H(t))\right)$
@@
with the convention that $\mathrm{rank}\,P_{(\lambda_{1},\lambda_{2})}(A)=-\mathrm{rank}\,P_{(\lambda_{2},\lambda_{1})}(A)$ if $\lambda_{2}<\lambda_{1}$. To use this formula of the spectral flow, we used here that
@@
which comes from the fact that $H(1)=\Sigma\,\widetilde{\oplus}\,H(0)$.
@@
$\mathrm{Ran}\left(P_{(E+a_{i},E+a_{i+1})}\left(H(t_{i})\right)\right)=\mathrm{Ran}\left(P_{(E+a_{i},E+a_{i+1})}\left(H_{\ell}(t_{i})\right)\right)\text{ for}i=1,\cdots,M-1$

### v3

@@ 3.2. Dislocated Jacobi operators
In order to study the right–half Jacobi operator $H^{\Sigma}_{R}(t)$, we study yet another dislocated model. The idea to study this model comes from Hempel and Kohlmann in [24, 23, 25], and was used in [18, 20] to compute spectral flows for Schrödinger operators acting on $L^{2}(\mathbb{R})$. In this setting, the dislocated model is modelled by a Schrödinger operator of the form $H_{t}:=-\partial_{xx}^{2}+V_{t}(x)$ with the dislocated potential
@@
where $V$ is $1$–periodic. The map $t\mapsto H_{t}$ is norm–resolvent continuous and $1$-periodic, and we recover periodic models at the endpoints. As mentioned in the introduction, a difficulty with the discrete model is that it is not possible to create such a continuous shift. Our workaround is to introduce a translation equivariant dislocated model $\widetilde{H}^{\Sigma}(t)$.
@@
See also Figure 1.3. The operator $\widetilde{H}^{\Sigma}(0)$ is exactly the periodic Jacobi operator $H$, and its spectrum is $\sigma_{\rm{bulk}}$. For the operator $\widetilde{H}^{\Sigma}(1)$, we note the following. Let us write ={0}($\mathbb{Z}=\{0\}\cup(\mathbb{Z}\setminus\{0\})$, and $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})=\mathbb{C}^{N}\oplus\ell^{2}(\mathbb{Z}\setminus\{0\},\mathbb{C}^{N})$. With respect to this decomposition, we write
@@
$\widetilde{H}^{\Sigma}(1)=\Sigma\,\widetilde{\oplus}\,\widetilde{H}^{\Sigma}(0),$ (14)
@@ 3.3. Computation of the spectral flow for the dislocated model
This result was first proved in [24, 23, 25] in the case of Schrödinger operators acting on the continuum. The proof presented here is similar, but more complete, and uses solely tools from spectral theory.
@@ 3.3.1. The finite periodic dislocated model
At $t=1$, apart from the presence $\Sigma$ in the middle, we also recognize a periodic model, and we have a relation of the form (compare with (14))
$\widetilde{H}_{\ell}^{\Sigma}(1)=\Sigma\,\widetilde{\oplus}\,H_{\ell-1},\qquad\text{hence}\qquad\sigma\left(\widetilde{H}_{\ell}^{\Sigma}(1)\right)=\sigma\left(H_{\ell-1}\right)\bigcup\{\Sigma\},$
@@ 3.3.2. From the finite dislocated model to the infinite dislocated model
In the previous section, we proved that the spectral flow of the supercell dislocated model $\widetilde{H}^{\Sigma}_{\ell}$ is $-{\mathcal{N}}(E)$ for all $\ell\geq 5$. In this section, we justify that one may take the limit $\ell\to\infty$ and deduce that spectral flow is also $-{\mathcal{N}}(E)$ for the full dislocated model $\widetilde{H}^{\Sigma}$. Throughout this subsection, we drop the superscript $\Sigma$, and the tilde notation for clarity: we will simply write $H(t)$ and $H_{\ell}(t)$ for the dislocated models acting on $\ell^{2}(\mathbb{Z},\mathbb{C}^{N})$ and $\ell^{2}(\Omega_{\ell},\mathbb{C}^{N})$ respectively. Recall that at $t=0$, the operators $H(0)$ and $H_{\ell}(0)$ are periodic, and that
@@
$\chi(x)=1\quad\forall x\in[-1/4,1/4],\qquad\chi(x)=0\quad\forall x\in\mathbb{R}\setminus(-1/3,1/3),$
and define $\chi_{\ell}$ as the multiplication operator by $\chi(\cdot/\ell)$ in ${\mathcal{H}}_{\ell}$. Note that this function is supported in $[-\ell/3,\ell/3]\subset\Omega_{\ell}$. We will use the following identities throughout this section.
@@
    $H(t)i_{\ell}\chi_{\ell}=i_{\ell}H_{\ell}(t)\chi_{\ell}$ from ${\mathcal{H}}_{\ell}$ to ${\mathcal{H}}$, and $\chi_{\ell}i_{\ell}^{*}H(t)=\chi_{\ell}H_{\ell}(t)i_{\ell}^{*}$ from ${\mathcal{H}}$ to ${\mathcal{H}}_{\ell}$,
@@
    For any multiplication operator $\phi$ on ${\mathcal{H}}_{\ell}$ with support in $\Omega_{\ell}\setminus[-\ell/4,\ell/4]$, we have $H_{\ell}(t)\phi=H_{\ell}(0)\phi$ as operators from ${\mathcal{H}}_{\ell}$ to ${\mathcal{H}}_{\ell}$.
    Similarly, For any multiplication operator $\widetilde{\phi}$ on ${\mathcal{H}}$ with support in $\mathbb{Z}\setminus[-\ell/4,\ell/4]$, we have $H(t)\widetilde{\phi}=H(0)\widetilde{\phi}$ as operators from ${\mathcal{H}}$ to ${\mathcal{H}}$.
@@
    The operator $R_{\ell}:=\left[H_{\ell},\chi_{\ell}\right]=H_{\ell}(t)\chi_{\ell}-\chi_{\ell}H_{\ell}(t)$ acting on ${\mathcal{H}}_{\ell}$ is independent of $t$, and satisfies $\|R_{\ell}\|_{{\rm{op}},\ell}\leq\frac{C}{\ell}$.
@@
For the first point, we note that for all $\Psi\in{\mathcal{H}}_{\ell}$, the function $\chi_{\ell}\Psi$ has support in $[-\ell/3,\ell/3]$. Whenever $\ell\geq 9$, this implies that $H_{\ell}(t)\chi_{\ell}\Psi$ has support in $[-\ell/3-1,\ell/3+1]$, hence $H_{\ell}(t)\chi_{\ell}\Psi=H(t)i_{\ell}\chi_{\ell}\Psi$. The second relation is the adjoint of the first.
For the second point, we note that $H_{\ell}(t)=H_{\ell}(0)+M_{\ell}(t)$, where $M_{\ell}(t):=i_{\ell}M(t)i_{\ell}^{*}$, and where $M(t)$ is a finite rank operator with nonzero matrix elements only in the entries $\{-1,0,1\}$. So $M_{\ell}(t)\phi=0$ whenever $\ell>4$, since we assumed that $\phi=0$ in $\Omega_{\ell}\cap[-\ell/4,\ell/4]\supset\{-1,0,1\}$. The argument for $\widetilde{\phi}$ is identical.
For the third point, since $\chi_{\ell}$ is equal to the identity in the support of $M$ for all $\ell>8$, we have $M_{\ell}(t)\chi_{\ell}=\chi_{\ell}M_{\ell}(t)$. In particular, $[H_{\ell}(t),\chi_{\ell}]=[H_{\ell}(0),\chi_{\ell}]$ is independent of $t$. Next, we find that
$\left[H_{\ell}(0)\chi_{\ell}-\chi_{\ell}H_{\ell}(0)\right]_{ij}=\begin{cases}a\ (\chi_{\ell}(n+1)-\chi_{\ell}(n))&\quad\text{if}\quad(i,j)=(n,n+1)\\ a^{*}(\chi_{\ell}(n-1)-\chi_{\ell}(n))&\quad\text{if}\quad(i,j)=(n,n-1)\\ 0&\quad\text{else}\end{cases}.$
@@
$R_{\ell}:=H_{\ell}(t)\chi_{\ell}-\chi_{\ell}H_{\ell}(t)=H_{\ell}(0)\chi_{\ell}-\chi_{\ell}H_{\ell}(0)\quad\text{satisfies}\quad\|R_{\ell}\|_{{\rm{op}},\ell}\leq 2\|a\|_{\rm op}\frac{C}{\ell}.\qed$
First, we prove that any complex number in the resolvent set of $H(t)$ is in the resolvent set of $H_{\ell}(t)$ for sufficiently large values of $\ell$.
@@
Since the essential spectrum of $H(t)$ coincides with $\sigma_{\rm{bulk}}$, we find that $\operatorname{dist}(\lambda,\sigma_{\rm{bulk}})\geq\varepsilon$. Recall that $\sigma(H_{\ell}(0))\subset\sigma_{\rm{bulk}}$, so in particular $H_{\ell}(0)-\lambda$ is invertible with $\|(H_{\ell}(0)-\lambda)^{-1}\|_{\rm op}\leq\varepsilon^{-1}$. The same bound holds for $(H(t)-\lambda)^{-1}$. We combine both resolvents to construct an approximate inverse for $H_{\ell}(t)-\lambda$. Indeed, for any $\ell\geq 9$, the identities from Lemma 3.6 give
$\left(H_{\ell}(t)-z\right)\chi_{\ell}i_{\ell}^{*}=\chi_{\ell}\left(H_{\ell}(t)-z\right)i_{\ell}^{*}+R_{\ell}i_{\ell}^{*}=\chi_{\ell}i_{\ell}^{*}\left(H(t)-z\right)+R_{\ell}i_{\ell}^{*}$ (15)
as operators from ${\mathcal{H}}$ to ${\mathcal{H}}_{\ell}$. Together with the fact that $i_{\ell}^{*}i_{\ell}={\mathds{1}}_{\ell}:={\mathds{1}}_{{\mathcal{H}}_{\ell}}$, we get the following identity, as operators from ${\mathcal{H}}_{\ell}$ to ${\mathcal{H}}_{\ell}$,
$(H_{\ell}(t)-\lambda)\left\{\chi_{\ell}i_{\ell}^{*}(H(t)-\lambda)^{-1}i_{\ell}+({\mathds{1}}_{\ell}-\chi_{\ell})(H_{\ell}(0)-\lambda)^{-1}\right\}$
$\quad=\chi_{\ell}i_{\ell}^{*}i_{\ell}+R_{\ell}i_{\ell}^{*}(H(t)-\lambda)^{-1}i_{\ell}+({\mathds{1}}_{\ell}-\chi_{\ell})-R_{\ell}(H_{\ell}(0)-\lambda)^{-1}$
$\quad={\mathds{1}}_{\ell}+R_{\ell}\left\{i_{\ell}^{*}(H(t)-\lambda)^{-1}i_{\ell}-(H_{\ell}(0)-\lambda)^{-1}\right\}:={\mathds{1}}_{\ell}+\mathcal{R}_{\ell}(t).$
@@
$\left\|\mathcal{R}_{\ell}(t)\right\|_{{\rm{op}},\ell}\leq\frac{2C}{\varepsilon\ell},$
@@
$\|(H_{\ell}(t)-\lambda)^{-1}\|_{{\rm{op}},\ell}$ $\leq\left\|\chi_{\ell}i_{\ell}^{*}(H(t)-\lambda)^{-1}i_{\ell}+({\mathds{1}}_{\ell}-\chi_{\ell})(H_{\ell}(0)-\lambda)^{-1}\right\|_{{\rm{op}},\ell}\,\|({\mathds{1}}_{\ell}+\mathcal{R}_{\ell}(t))^{-1}\|_{{\rm{op}},\ell}$
@@
$\forall z\in{\mathscr{C}},\quad\|(z-H_{\ell}(t))^{-1}\|_{\rm op}\leq\frac{4}{\varepsilon},\quad\text{and}\quad\|(z-H(t))^{-1}\|_{\rm op}\leq\frac{1}{\varepsilon}.$
We use the localization function $\chi_{\ell}$ and $\phi_{\ell}:=\sqrt{1-\chi_{\ell}^{2}}$, such that $\chi_{\ell}^{2}+\phi_{\ell}^{2}=1$. The support of $\chi_{\ell}$ is in $[-\ell/3,\ell/3]$, the support of $\phi_{\ell}$ is in the complement of $(-\ell/4,\ell/4)$. The third point of Lemma 3.6 (for $\chi_{\ell}$, but the proof is similar for $\phi_{\ell}$) shows that there is a constant $C$ so that
$R_{\ell}:=[H_{\ell}(t),\chi_{\ell}]\quad\text{and}\quad R_{\ell}^{\phi}:=[H_{\ell}(t),\phi_{\ell}]\quad\text{satisfy}\quad\left\|R_{\ell}\right\|_{{\rm{op}},\ell}\leq\frac{C}{\ell},\quad\left\|R_{\ell}^{\phi}\right\|_{{\rm{op}},\ell}\leq\frac{C}{\ell}.$ (16)
Consider $z\in{\mathscr{C}}$, and recall (15) which states that
$\chi_{\ell}i_{\ell}^{*}(H(t)-z)=(H_{\ell}(t)-z)\chi_{\ell}i_{\ell}^{*}-R_{\ell}.$
Composing this identity with $(H_{\ell}(t)-z)^{-1}$ on the left and $(H(t)-z)^{-1}i_{\ell}\chi_{\ell}$ on the right gives
$(H_{\ell}(t)-z)^{-1}\chi_{\ell}^{2}=\chi_{\ell}i_{\ell}^{*}(H(t)-z)^{-1}i_{\ell}\chi_{\ell}+{\mathcal{R}}^{(1)}_{\ell}(z),$ (17)
with $\|{\mathcal{R}}^{(1)}_{\ell}(z)\|_{{\rm{op}},\ell}\leq\frac{8C}{\varepsilon^{2}\ell}$. On the other hand, we have $H_{\ell}(t)\phi_{\ell}=H_{\ell}(0)\phi_{\ell}$ as operators on ${\mathcal{H}}_{\ell}$ (see Lemma 3.6(2)), hence, taking adjoints, $\phi_{\ell}H_{\ell}(t)=\phi_{\ell}H_{\ell}(0)$. Writing
$\phi_{\ell}(H_{\ell}(0)-z)=\phi_{\ell}(H_{\ell}(t)-z)=(H_{\ell}(t)-z)\phi_{\ell}-R_{\ell}^{\phi}$
and multiplying by $(H_{\ell}(t)-z)^{-1}$ on the left and $(H_{\ell}(0)-z)^{-1}\phi_{\ell}$ on the right gives
$(H_{\ell}(t)-z)^{-1}\phi_{\ell}^{2}=\phi_{\ell}(H_{\ell}(0)-z)^{-1}\phi_{\ell}+{\mathcal{R}}^{(2)}_{\ell}(z),$ (18)
again with $\|{\mathcal{R}}^{(2)}_{\ell}(z)\|_{{\rm{op}}}\leq\frac{8C}{\varepsilon^{2}\ell}$. Summing (17) and (18) gives
@@
$P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t))=\chi_{\ell}i_{\ell}^{*}P_{(\lambda_{1},\lambda_{2})}(H(t))i_{\ell}\chi_{\ell}+\phi_{\ell}P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(0))\phi_{\ell}+\frac{1}{2{\mathrm{i}}\pi}\oint_{\mathscr{C}}({\mathcal{R}}_{\ell}^{(1)}+{\mathcal{R}}_{\ell}^{(2)})(z){\mathrm{d}}z.$
Since $(\lambda_{1},\lambda_{2})$ lies in a gap of $\sigma_{\rm{bulk}}$, the second term is actually zero. The last term can be bounded by $4C|{\mathscr{C}}|/(\pi\varepsilon^{2}\ell)$. Thus, we can find $\ell_{0}$ sufficiently large such that for all $\ell\geq\ell_{0}$, we have
$\left\|P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t))-\chi_{\ell}i_{\ell}^{*}P_{(\lambda_{1},\lambda_{2})}(H(t))i_{\ell}\chi_{\ell}\right\|_{\rm{op}}<1.$
@@
For the opposite inequality, we have to define the cut-off function $\widetilde{\phi}_{\ell}$ such that $\widetilde{\phi}_{\ell}^{2}+i_{\ell}\chi_{\ell}^{2}i^{*}_{\ell}={\mathds{1}}_{\mathcal{H}}$. Multiplying (15) on the left by $i_{\ell}\chi_{\ell}(H_{\ell}(t)-z)^{-1}$ and on the right by $(H(t)-z)^{-1}$ gives the identity
$i_{\ell}\chi_{\ell}^{2}i_{\ell}^{*}(H(t)-z)^{-1}=i_{\ell}\chi_{\ell}(H_{\ell}(t)-z)^{-1}\chi_{\ell}i_{\ell}^{*}+\widetilde{{\mathcal{R}}}^{(1)}_{\ell}(z),$
with ~$\|\widetilde{{\mathcal{R}}}^{(1)}_{\ell}(z)\|_{{\rm{op}},\ell}\leq\frac{8C}{\varepsilon^{2}\ell}$. The analogue of (18) reads
$\widetilde{\phi}_{\ell}^{2}(H(t)-z)^{-1}=\widetilde{\phi}_{\ell}(H(0)-z)^{-1}\widetilde{\phi}_{\ell}+\widetilde{{\mathcal{R}}}^{(2)}_{\ell}(z).$
with ~$\|\widetilde{{\mathcal{R}}}^{(2)}_{\ell}(z)\|_{{\rm{op}},\ell}\leq\frac{8C}{\varepsilon^{2}\ell}$. We then apply the Cauchy formula to
$(H(t)-z)^{-1}$ $=i_{\ell}\chi_{\ell}(H_{\ell}(t)-z)^{-1}\chi_{\ell}i_{\ell}^{*}+\widetilde{\phi}_{\ell}(H(0)-z)^{-1}\widetilde{\phi}_{\ell}+(\widetilde{{\mathcal{R}}}_{\ell}^{(1)}+\widetilde{{\mathcal{R}}}_{\ell}^{(2)})(z),$
and note that the second term vanishes as before. Thus, Lemma 3.8 gives for all sufficiently large values of $\ell$
$\mathrm{rank}(P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t)))\geq\mathrm{rank}\left(i_{\ell}\chi_{\ell}P_{(\lambda_{1},\lambda_{2})}(H_{\ell}(t))\chi_{\ell}i_{\ell}^{*}\right)\geq\mathrm{rank}\left(P_{(\lambda_{1},\lambda_{2})}(H(t))\right),$
@@
with the convention that $\mathrm{rank}\,P_{(\lambda_{1},\lambda_{2})}(A)=-\mathrm{rank}\,P_{(\lambda_{2},\lambda_{1})}(A)$ if $\lambda_{2}<\lambda_{1}$. We used here that
@@
which comes from the fact that $H(1)=\Sigma\,\widetilde{\oplus}\,H(0)$, so that there are no boundary terms
@@
$\mathrm{Ran}\left(P_{(E+a_{i},E+a_{i+1})}\left(H(t_{i})\right)\right)=\mathrm{Ran}\left(P_{(E+a_{i},E+a_{i+1})}\left(H_{\ell}(t_{i})\right)\right)\text{ for}i=1,\cdots,M-1,$